If a a paint mixture uses 60 pounds of solvent for 75 gallons of paint, how many pounds of solvent are needed for 90 gallons of paint?
Answer by checkley71(8403)

60/75=x/90 cross multiply
75x=60*90
75x=5400
x=5400/75
x=72 pound of solvent are required.